# Automate Strategy Finding with LLM in Quant investment

### v1

@@ 5.4. Discussion
The term “formulaic alpha” is often associated with investing, the concept of simple and interpretable formulaic predictors, which can be combined into more expressive models and many more application scenarios. The concept of the “formulaic alpha” can solve many forecasting problems in other subjects such as energy consumption prediction, anomaly detection, and biomedical applications. (deb2017review, ; ashok2018online, ; tuarob2017feeling, )The work also has potential future work, The current framework utilizes a multi-Agent architecture to evaluate market conditions and select alpha factors. Future work can explore transitioning this architecture to a Mixture of Experts (MoE) structure, which can improve learning efficiency and model performance. MoE can dynamically allocate computational resources to different ”experts” based on the specific characteristics of the input data, thereby enhancing the model’s ability to learn and distinguish category seed alphas features more effectively. (masoudnia2014mixture, )Another promising avenue for future research is the incorporation of financial knowledge graph techniques into the strategy development process. Financial markets are inherently complex and interconnected, with numerous dependencies between stocks, sectors, and economic indicators. (wang2023methods, ) By introducing the knowledge graph information may help in our stock selection phase and achieve better performance in the future.

### v2

@@ 5.4. Discussion
The term “formulaic alpha” is often associated with investing, the concept of simple and interpretable formulaic predictors, which can be combined into more expressive models and many more application scenarios. The concept of the “formulaic alpha” can solve many forecasting problems in other subjects such as energy consumption prediction, anomaly detection, and biomedical applications (deb2017review, ; ashok2018online, ; tuarob2017feeling, ). The work also has potential future work, The current framework utilizes a multi-Agent architecture to evaluate market conditions and select alpha factors. Future work can explore transitioning this architecture to a Mixture of Experts (MoE) structure, which can improve learning efficiency and model performance. MoE can dynamically allocate computational resources to different ”experts” based on the specific characteristics of the input data, thereby enhancing the model’s ability to learn and distinguish category seed alphas features more effectively (masoudnia2014mixture, ). Another promising avenue for future research is the incorporation of financial knowledge graph techniques into the strategy development process. Financial markets are inherently complex and interconnected, with numerous dependencies between stocks, sectors, and economic indicators (wang2023methods, ). By introducing the knowledge graph information may help in our stock selection phase and achieve better performance in the future.